# Robustness in Large Language Models: A Survey of Mitigation Strategies and Evaluation Metrics

### v1

@@ 4.2 In-processing Strategies
Adversarial Training (AT) strengthens models by intentionally injecting adversarial examples into their training data, thereby forcing the model to learn more robust features. It can be formulated as a minimax optimization problem: the inner loop maximizes the loss by finding the worst-case adversarial example within a defined perturbation budget, while the outer loop minimizes the loss on these adversarial examples (along with clean examples) to update the model parameters (Bai et al., 2021a).
Figure 4: The robustness of Large Language Models (LLMs) can be systematically fortified through a layered pipeline. Beginning with data collection and annotation, diverse datasets and rigorous labelling establish a foundational resilience. Pre-processing strategies then clean, augment, and balance the data to mitigate biases and noise. During training and fine-tuning, in-processing techniques like adversarial learning and alignment strategies address vulnerabilities directly. As part of inference, intra-processing strategies dynamically adjust inputs (e.g., sanitising adversarial queries) to counter real-time risks. Finally, post-processing safeguards act as a last layer, refining outputs via alignment checks and ethical filters to ensure reliability. Together, these stages create a unified defence, ensuring adaptability and resilience across the LLM lifecycle.
Early approaches involved adapting methods from other domains like generating attacks via word substitutions guided by gradient information or word importance scores (Liu & Sun, 2023; Weng, 2023) or performing perturbations directly in the continuous embedding space and then mapping these perturbed embeddings back to discrete tokens, often using a masked language model (MLM) head (Altinisik et al., 2023). Aiming to improve robustness generalization by learning a distribution over adversarial examples around each clean data point, rather than finding a single worst-case point estimate, Dong et al. (2021) proposes Adversarial Distributional Training (ADT), which uses an entropy regularize. Further, Xhonneux et al. (2024) proposes a fast adversarial training algorithm C-AdvUL (Continuous-Adversarial Unlikelihood) and CAPO (Continuous-Adversarial IPO). C-AdvUL combines an unlikelihood loss (penalizing harmful outputs, rewarding safe ones) on perturbed inputs with a standard utility loss on clean data. CAPO integrates continuous attacks with Identity Preference Optimisation (IPO), a DPO variant (Rafailov et al., 2024), using the IPO loss on perturbed inputs and relying on the KL divergence term inherent in IPO to maintain utility without requiring a separate utility dataset. Distributionally Robust Optimization (DRO) (Rahimian & Mehrotra, 2022; Kuhn et al., 2025) is another work for optimizing model performance under worst-case conditions over a specified distribution or ambiguous set of inputs, rather than just point-wise adversarial examples. Recent work like PEARL (Permutation-resilient learning) and Dr. DPO (Distributionally Robustifying DPO) has also applied these principles to specific LLM robustness challenges (Wu et al., 2025; Chen et al., 2025a).
@@
However, scaling AT to LLMs also introduces several challenges. First, the discrete nature of text complicates gradient-based perturbation strategies, where small token-level modifications must preserve semantic coherence, and gradients derived from differentiable surrogates (e.g., word embeddings) do not always map meaningfully to discrete token substitutions. Second, the computational overhead grows prohibitive: generating adversarial examples via iterative methods (e.g., Projected Gradient Descent) or combinatorial search over token sequences is resource-intensive, and fine-tuning massive LLMs on such augmented datasets amplifies training costs exponentially. Third, AT often induces a robustness-utility trade-off, where improvements in adversarial accuracy come at the expense of degraded performance on clean data, necessitating careful balancing. Finally, robustness gains are frequently attack-specific as models trained against one perturbation type (e.g., synonym swaps) may remain vulnerable to unseen attack strategies (e.g., paraphrasing or structural perturbations). And so efficiency and integration with alignment objectives are becoming central themes, with continuous attacks and DRO (Dong et al., 2021; Chen et al., 2025a) emerging as promising directions.
Regularization
Regularization methods add constraints or penalties to the training objective to prevent overfitting and encourage desirable model properties, including robustness. A key approach to this is Adversarial Contrastive Learning (ACL) (Jiang et al., 2020), which integrates robustness directly into representation learning by aligning embeddings of clean, augmented, and adversarially perturbed inputs. Recent advances, such as robustness-aware coreset selection and adversarial invariant regularization terms (Xu et al., 2023), further streamline ACL by focusing training on critical data subsets and explicitly penalizing sensitivity to adversarial distortions. Mahabadi et al. (2020) developed Debiased Focal Loss (DBL), which adaptively down-weights biased examples identified by an auxiliary bias-only model. Complementary work by Utama et al. (2020) introduced confidence regularization to balance in-distribution accuracy and out-of-distribution generalization, specifically preventing overconfidence in biased examples without sacrificing overall learning capacity. Further, Garimella et al. (2021) proposed another regularization approach that identifies potentially biased terms through lexical co-occurrence analysis, assigning bias scores to adjectives and adverbs ($W$), which are then incorporated into the model’s optimization process. Several studies have also been conducted to obtain a fair model by using Regularization (Ravichandran et al., 2020; Xu et al., 2025; Heck et al., 2024; Mou et al., 2024; Li et al., 2022). Together, these methods demonstrate how structured regularization can balance robustness, efficiency, and generalization in modern language models.
@@
Aligning LLMs with human values during training aims to ensure these systems are helpful and reliable. Techniques like supervised fine-tuning (SFT) (Sun, 2024; Luo et al., 2024) on high-quality data and reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Christiano et al., 2023), which uses human preferences to shape model behaviour, are widely adopted. Newer methods, such as Direct Preference Optimization (DPO), simplify alignment by training models directly on preference-ranked outputs, while variants like ORPO, KTO, and AlphaPO (Hong et al., 2024; Ethayarajh et al., 2024; Gupta et al., 2025) seek to enhance efficiency or performance. While primarily focused on safety and helpfulness, alignment inherently contributes to robustness against certain failure modes, particularly attacks designed to elicit harmful or biased content. For example, approaches like Safe RLHF (Tran et al., 2025) integrate safety-specific rewards to balance helpfulness with risk mitigation.
@@
However, alignment introduces its own challenges. First, models may become overly sensitive to prompt phrasing, leading to inconsistent responses or unnecessary refusals of benign queries (Oh & Demberg, 2025; Tran et al., 2025). Second, alignment might overlook emerging vulnerabilities, as pre-alignment evaluations often fail to predict post-alignment weaknesses. Third, reward hacking can occur, where models exploit flaws in their training objectives to maximize scores without genuinely adhering to human intent (Herrera-Poyatos et al., 2025). Finally, methods like RLHF demand extensive computational resources and rely heavily on high-quality, diverse training data, limiting scalability. Addressing these trade-offs remains critical to developing reliable, ethically aligned AI systems.
@@
Acknowledging these challenges, researchers are actively developing strategies to strengthen the alignment process and mitigate its inherent risks. One approach integrates AT with alignment objectives, such as combining continuous adversarial perturbations with preference optimization frameworks (Xhonneux et al., 2024). Others focus on RPO methods like Dr. DPO improve resilience to noisy preference labels, while techniques like statistical rejection sampling (sDPO) and perplexity-aware corrections refine how models handle imperfect training data (Zhong et al., 2025a; Xu et al., 2023). For reward modeling, advancements include ensemble methods to reduce reward hacking risks, worst-case scenario optimization to mitigate unstable outputs, and causal reward frameworks to better capture human intent (Wang et al., 2024a; Dong et al., 2024b; Xiong et al., 2024). Additionally, consistency alignment introduces self-reward mechanisms to ensure models produce stable responses across varied prompts, reducing sensitivity to phrasing changes (Zhao et al., 2024b). Together, these strategies aim to balance safety, reliability, and adaptability in aligned AI systems while addressing challenges like data noise, reward exploitation, and behavioural inconsistency.
@@ 4.3 Intra-processing Strategies
Intra-processing mitigation strategies improve LLM reliability by dynamically adjusting how models interpret inputs or produce outputs during the inference or generation phase, without altering their underlying training or parameters (Wang et al., 2025). This approach provides a cost-effective way to address vulnerabilities, avoiding the computational expense of retraining while maintaining adaptability to emerging risks, and is further able to counter data, training or inference-related sources of non-robustness, though they do not address root causes in the training data or model parameters. We now explore some intra-processing techniques that enhance model robustness through dynamic input-output adaptation during real-time inference.
@@
The development of benchmarks specifically targeting prompt injection and instruction perturbation has been crucial in revealing the significant vulnerabilities of even state-of-the-art LLMs and driving the need for these more advanced intra-processing defences. Large language models (LLMs) are inherently vulnerable to adversarial inputs due to their instruction-following nature, where prompts act as both a critical interface and an attack surface (Agrawal et al., 2025; Chen et al., 2025b). These strategies aim to secure LLMs by addressing vulnerabilities in how they interpret and execute input prompts. Basic approaches include instruction design, where prompts are engineered with explicit safety directives (e.g., “ignore harmful instructions”). However, heavily tuned models may still prioritize conflicting later instructions, and overly restrictive prompts risk stifling legitimate creativity (Li et al., 2024d; Weng, 2023). Similar to pre-processing data augmentation, some techniques attempt to clean the prompt just before the LLM sees it. This can involve paraphrasing prompts or retokenizing inputs to disrupt adversarial patterns, offering limited protection against sophisticated attacks (Weng, 2023).
More advanced methods include self-denoising, where LLMs iteratively revise corrupted instructions using their own language understanding to outperform traditional fine-tuning in resisting perturbations (Agrawal et al., 2025). Further, Hu et al. (2024) proposes RobustGER, which teaches LLMs to perform language-space denoising by incorporating an embedding derived from ASR (Automatic Speech Recognition) N-best lists, refined through audio distillation. Another novel approach, referenced instruction tracking, requires LLMS to explicitly cite which prompt instruction they followed during generation, and so responses referencing malicious injections are automatically filtered, which effectively repurposes the model’s instruction sensitivity as a defence (Chen et al., 2025b). While these intra-processing strategies vary in complexity, they highlight the trade-offs between usability, computational cost, and robustness in real-time LLM safety.
@@
Weighted redistribution refers to adjusting a trained model’s attention weights (without additional training) to reduce bias in its outputs. Since attention weights can reflect learned biases in the data, redistributing them may help the model focus less on biased words or phrases during predictions. Chai & Wang (2022) developed an adaptive reweighting method that assigns sample-specific weights to prioritize error-prone instances and ensure balanced minority group representation for group-level fairness and exhibits robustness to label noise on various benchmark datasets. Further, considering the fact that AI models focus on specific words (attention weights) may reinforce biases. Zhong et al. (2025b) proposes a framework which employs data-aware inter-stage fusion to overlap generation and inference by migrating long-tailed samples and model-aware intra-stage fusion with a fused pipeline schedule to mitigate pipeline bubbles in the training stage to prepare the context of inference tasks.
@@
The methods that guide how LLMs select tokens to generate text also play a subtle but important role in balancing output quality and robustness. While standard approaches like greedy decoding (choosing the most probable token) or beam search are widely used, they often struggle under adversarial conditions or when generating diverse, reliable outputs (Beyer et al., 2025). Despite their limitations, research has largely prioritized input- or output-level defences (e.g., prompt engineering, response filtering) over reimagining decoding itself as a primary robustness mechanism. This gap persists even as adversarial attacks exploit decoding weaknesses, such as using greedy search or Greedy Coordinate Gradient (GCG) (Kumar et al., 2025a) to craft optimal input perturbations. The relative neglect of defensive decoding strategies may stem from practical hurdles, such as the computational complexity of altering core generation processes compared to simpler input/output interventions.
Emerging approaches aim to address these gaps by refining how LLMs navigate token selection. Uncertainty-aware decoding, for instance, could suppress generations when the model’s confidence is low, reducing errors in high-stakes scenarios. Ensemble decoding combines outputs from varied model states or prompts to dilute adversarial influences, while constrained decoding enforces safety rules (e.g., blocking harmful phrases) during token selection (Shi et al., 2024b; Gu et al., 2024; Daheim et al., 2025). Son et al. (2025) formalizes this challenge as a maximin game, introducing RMOD (Robust Multi-Objective Decoding), a method aimed at maximizing worst-case rewards through Nash equilibrium optimization. RMOD achieves this by balancing competing objectives, specifically, identifying a Nash equilibrium between reward weight allocations and the sampling policy during text generation.
@@
Beyond prompt manipulation and decoding, other intra-processing techniques involve adapting the model or its execution environment at inference time. While efficient inference infrastructure does not directly enhance robustness, it enables the deployment of adaptive systems that can improve real-world reliability (Manvi et al., 2024). For instance, LeVine et al. (2024) evaluates the performance and calibration of reward models under distribution shift and finds that accuracy degrades significantly, particularly for OOD responses. Some of key strategies include test-time adaptation (TTA), which makes minor, instance-specific adjustments to improve handling of rare or unfamiliar data, such as medical cases outside training distributions (Snell et al., 2024). Architectural interventions explore structural modifications during inference, like removing or swapping transformer layers, revealing that models retain most functionality despite such changes, which is a sign of inherent redundancy and fault tolerance (Lad et al., 2024). Other work examines components like layer normalization, showing how architectural choices impact stability under perturbations (Jha & Reagen, 2024). Finally, efficient inference engines optimize computational bottlenecks (e.g., attention mechanisms) using memory optimization and adaptive computation, enabling complex real-time processes like retrieval-augmented generation while balancing speed and accuracy (Ye et al., 2025). Together, these approaches highlight how real-time adaptability and structural insights can bolster LLM reliability without sacrificing efficiency.

### v2

@@ 4.2 In-processing Strategies
Adversarial Training (AT) strengthens models by intentionally injecting adversarial examples into their training data, thereby forcing the model to learn more robust features. It can be formulated as a minimax optimization problem: the inner loop maximizes the loss by finding the worst-case adversarial example within a defined perturbation budget, while the outer loop minimizes the loss on these adversarial examples (along with clean examples) to update the model parameters (bai2021recentadvancesadversarialtraining).
Early approaches involved adapting methods from other domains like generating attacks via word substitutions guided by gradient information or word importance scores (liu-sun-2023-end; weng2023attack) or performing perturbations directly in the continuous embedding space and then mapping these perturbed embeddings back to discrete tokens, often using a masked language model (MLM) head (altinisik-etal-2023-impact). While some methods also show that AT can improve both generalization and robustness for a wide range of NLP tasks. liu2020adversarialtraininglargeneural introduce ALUM (Adversarial training for large neural LangUage Models), a method designed for both pre-training and fine-tuning of Transformer-based language models. It regularizes the training objective by applying perturbations in the embedding space that maximize adversarial loss, effectively enforcing label smoothness in local neighborhoods. The approach builds on virtual adversarial training, with the objective defined as:
$\min_{\theta}E_{(x,y)\sim D}\left[l(f(x;\theta),y)+\alpha\max_{\delta}l(f(x+\delta;\theta),f(x;\theta))\right]$
Here, $f(x;\theta)$ represents the neural language model parameterized by $\theta$, $l(\cdot,\cdot)$ is the loss function (e.g., cross-entropy), $(x,y)$ is an input-output pair from the training dataset $D$, and $\delta$ is a small perturbation applied to the embedding $x$. The term $\alpha$ is a hyperparameter that controls the trade-off between the standard empirical risk and the adversarial regularization term. Further, xhonneux2024efficientadversarialtrainingllms propose a fast adversarial training algorithm C-AdvUL (Continuous-Adversarial Unlikelihood) and CAPO (Continuous-Adversarial IPO). C-AdvUL applies the unlikelihood (UL) loss with continuous embedding attacks and incorporates an additional utility loss term, fine-tuning on a utility dataset $\mathcal{D}_{u}$, apart from the original dataset $\mathcal{D}$ :
$\min_{\theta}-\mathbb{E}_{(x,y,\hat{y})\in\mathcal{D}}\left[\log f_{\theta}(y|x+\delta(x,\hat{y}))-\log f_{\theta}(\hat{y}|x+\delta(x,\hat{y}))\right]-\mathbb{E}_{(x,y)\in\mathcal{D}_{u}}\left[\log f_{\theta}(y|x)\right]$
While, CAPO integrates continuous attacks with Identity Preference Optimisation (IPO), a DPO variant (rafailov2024directpreferenceoptimizationlanguage), using the IPO loss on perturbed inputs and relying on the KL divergence term inherent in IPO to maintain utility without requiring a separate utility dataset. The loss function for CAPO is:
$\min_{\theta}-\mathbb{E}_{(x,y,\hat{y})\in\mathcal{D}}\left[\ell_{\beta}\left(\log\frac{f_{\theta}(y|x+\delta(x,\hat{y}))}{f_{\theta_{0}}(y|x)}-\log\frac{f_{\theta}(\hat{y}|x+\delta(x,\hat{y}))}{f_{\theta_{0}}(\hat{y}|x)}\right)\right]$
[…]
Distributionally Robust Optimization (DRO) (Rahimian_2022; kuhn2025distributionallyrobustoptimization) is another work for optimizing model performance under worst-case conditions over a specified distribution or ambiguous set of inputs, rather than just point-wise adversarial examples. Recent work like PEARL (Permutation-resilient learning) and Dr. DPO (Distributionally Robustifying DPO) has also applied these principles to specific LLM robustness challenges (wu2025towards; chen2025pearlpermutationresilientllms).
@@
However, scaling AT to LLMs also introduces several challenges. First, the discrete nature of text complicates gradient-based perturbation strategies, where small token-level modifications must preserve semantic coherence, and gradients derived from differentiable surrogates (e.g., word embeddings) do not always map meaningfully to discrete token substitutions. Second, the computational overhead grows prohibitive: generating adversarial examples via iterative methods (e.g., Projected Gradient Descent) or combinatorial search over token sequences is resource-intensive, and fine-tuning massive LLMs on such augmented datasets amplifies training costs exponentially. Third, AT often induces a robustness-utility trade-off, where improvements in adversarial accuracy come at the expense of degraded performance on clean data, necessitating careful balancing. Finally, robustness gains are frequently attack-specific as models trained against one perturbation type (e.g., synonym swaps) may remain vulnerable to unseen attack strategies (e.g., paraphrasing or structural perturbations). And so efficiency and integration with alignment objectives are becoming central themes, with continuous attacks and DRO (dong-etal-2021-data; chen2025pearlpermutationresilientllms) emerging as promising directions.
Regularization and Constrained Optimization
Regularization methods add constraints or penalties to the training objective to prevent overfitting and encourage desirable model properties, including robustness. Recent advances highlight methods like BSR (Batch-wise Sum-to-Zero Regularization) to enforce zero-centered reward sums per batch, effectively constraining extreme reward magnitudes and mitigating hidden state norm dispersion (hong2025robustnessrewardmodelslanguage). It adds an additive regularization term to the Bradley-Terry (BT) loss in reward models: $L_{BT-BSR}=L_{BT}+\lambda\cdot L_{BSR}$. The BSR term is defined as
$L_{BSR}=\left(\frac{1}{2|B|}\sum_{i=1}^{|B|}\sum_{j\in\{w,l\}}r(x_{i},y_{i,j})\right)^{2}$
Here $|B|$ is the batch size and $\lambda$ is a weight hyperparameter. sharma2024information propose “information-guided regularization”, which integrates dropout with information-theoretic principles to guide fine-tuning, thereby improving generalization and robustness against overfitting. Similarly,vukadin2025largelanguagemodelsattribution propose Large Language Model Attribution Aligned Training (LAAT), which incorporates LLM-generated global task feature attributions into the training process of smaller networks through an attribution-matching regularization term, yielding superior performance in few-shot learning scenarios. The overall loss function in LAAT can be defined as:
$L(\theta)=\frac{1}{n}\sum_{i=1}^{n}\left(\ell_{BCE}(m_{\theta}(x_{i}),y_{i})+\gamma\ell_{MSE}\left(\frac{a(x_{i})}{\|a(x_{i})\|},\frac{s_{LLM}}{\|s_{LLM}\|}\right)\right)$
Here, $\ell_{BCE}$ is the standard binary cross-entropy loss, and $\ell_{MSE}$ is the mean squared error between the normalized attribution scores. The term $a(x_{i})$ represents the input gradient, $\nabla_{x_{i}}\ell_{BCE}(m_{\theta}(x_{i}),y_{i})$, which quantifies the sensitivity of the model’s loss with respect to each input feature $x_{i}$. $s_{LLM}$ is the vector of LLM-derived importance scores, and $\gamma$ is a weighting factor for the regularization term. The normalization by Euclidean norm, e.g., $\frac{a(x_{i})}{\|a(x_{i})\|}$, ensures that the regularization aligns the direction of feature importance rather than its raw magnitude. This regularization acts as an inductive bias, guiding the smaller model’s learning dynamics to conform to the LLM’s broader understanding of feature relevance. mahabadi2020endtoendbiasmitigationmodelling proposed Debiased Focal Loss (DBL) for NLU classification tasks, where biased training examples are down-weighted using an auxiliary bias-only model. While not developed for language model generation, the underlying idea of adaptive reweighting to reduce reliance on spurious correlations remains relevant. Likewise, utama2020mindtradeoffdebiasingnlu introduced confidence regularization to mitigate overconfidence on biased examples, striking a balance between in-distribution accuracy and OOD generalization. Similarly, (Garimella2021HeIV) designed a lexical co-occurrence based regularizer that assigns bias scores to adjectives and adverbs ($W$), incorporating them into the optimization process.
More broadly, regularization has also been explored for fairness and noisy-label robustness in machine learning (ravichandran2020fairxgboostfairnessawareclassificationxgboost; xu2025collapsedlanguagemodelspromote; heck2024learningnoisylabelsselftaught; Mou_Yue_Zhao_Wang_2024; li2022accuratefairnessimprovingindividual). While many of these works focus on classifiers rather than generative LMs, their principles in penalizing sensitivity to biased features, adjusting loss functions to handle noisy data, or promoting balanced performance across groups, are conceptually transferable to LLMs. Together, these methods demonstrate how structured regularization can be leveraged to balance robustness, fairness, efficiency, and generalization in modern language models.
@@
Aligning LLMs with human values during training aims to ensure these systems are helpful and reliable. (yao2025tokenconstraintdecodingimproves) develop Token Constraint Decoding (TCD), an inference-time algorithm that enforces alignment between token-level predictions to bolster robustness in noisy settings. Their extensive experiments reveal that TCD significantly restores performance degraded by input noise, yielding up to a 39% absolute gain for smaller models. Techniques like Supervised Fine-Tuning (SFT) (sun2024supervisedfinetuninginversereinforcement; luo2024robustftrobustsupervisedfinetuning) on high-quality data and Reinforcement Learning from Human Feedback (RLHF) (ouyang2022traininglanguagemodelsfollow; christiano2023deepreinforcementlearninghuman), which uses human preferences to shape model behaviour, are widely adopted. Newer RL methods, such as Direct Preference Optimization (DPO), simplify alignment by training models directly on preference-ranked outputs, while variants like ORPO (Odds Ratio Preference Optimization hong2024orpomonolithicpreferenceoptimization), KTO (ethayarajh2024ktomodelalignmentprospect), and AlphaPO (gupta2025alphaporewardshape) seek to enhance efficiency or performance. Inspired by DPO and KTO (safety) inroduce a novel constrained fine-tuning objective designed to protect the generative distribution of initial tokens. This formulation enhances the persistence of safety alignment under fine-tuning attacks while maintaining the model’s utility, but adapted to control the deviation from the initial generative distribution for each token position, similarly to the token-wise RL objectives in literature.
While primarily focused on safety and helpfulness, alignment inherently contributes to robustness against certain failure modes, particularly attacks designed to elicit harmful or biased content. For example, approaches like Safe RLHF (tran2025vulnerabilitymitigationsafetyalignedlanguage) integrate safety-specific rewards to balance helpfulness with risk mitigation.
@@
However, alignment introduces its own challenges. First, models may become overly sensitive to prompt phrasing, leading to inconsistent responses or unnecessary refusals of benign queries (Oh_Demberg_2025; tran2025vulnerabilitymitigationsafetyalignedlanguage). Second, alignment might overlook emerging vulnerabilities, as pre-alignment evaluations often fail to predict post-alignment weaknesses. Third, reward hacking can occur, where models exploit flaws in their training objectives to maximize scores without genuinely adhering to human intent (herrerapoyatos2025overviewmodeluncertaintyvariability). Finally, methods like RLHF demand extensive computational resources and rely heavily on high-quality, diverse training data, limiting scalability. Addressing these trade-offs remains critical to developing reliable, ethically aligned AI systems.
@@
Acknowledging these challenges, researchers are actively developing strategies to strengthen the alignment process and mitigate its inherent risks. One approach integrates AT with alignment objectives, such as combining continuous adversarial perturbations with preference optimization frameworks (xhonneux2024efficientadversarialtrainingllms). Others focus on RPO (Reward-aware Preference Optimization) methods like Dr. DPO (Distributionally Robustifying DPO wu2025towards) improve resilience to noisy preference labels, while techniques like Segment-Level DPO (SDPO kong2025sdposegmentleveldirectpreference) and perplexity-aware corrections refine how models handle imperfect training data (zhong2025comprehensivesurveyrewardmodels; xu2023efficientadversarialcontrastivelearning). For reward modeling, advancements include ensemble methods to reduce reward hacking risks, worst-case scenario optimization to mitigate unstable outputs, and causal reward frameworks to better capture human intent (ArmoRM; dong2024rlhf; xiong2024iterative). Additionally, consistency alignment introduces self-reward mechanisms to ensure models produce stable responses across varied prompts, reducing sensitivity to phrasing changes (zhao-etal-2024-improving). Together, these strategies aim to balance safety, reliability, and adaptability in aligned AI systems while addressing challenges like data noise, reward exploitation, and behavioural inconsistency.
@@ 4.3 Intra-processing Strategies
Intra-processing mitigation strategies improve LLM reliability by dynamically adjusting how models interpret inputs or produce outputs during the inference or generation phase, without altering their underlying training or parameters (wang2025surveyresponsiblellmsinherent). This approach provides a cost-effective way to address vulnerabilities, avoiding the computational expense of retraining while maintaining adaptability to emerging risks, and is further able to counter data, training or inference-related sources of non-robustness, though they do not address root causes in the training data or model parameters. We now explore some intra-processing techniques that enhance model robustness through dynamic input-output adaptation during real-time inference.
@@
The development of benchmarks specifically targeting prompt injection and instruction perturbation has been crucial in revealing the significant vulnerabilities of even state-of-the-art LLMs and driving the need for these more advanced intra-processing defences. LLMs are inherently vulnerable to adversarial inputs due to their instruction-following nature, where prompts act as both a critical interface and an attack surface (agrawal2025enhancingllmrobustnessperturbed; chen2025robustnessreferencingdefendingprompt). These strategies aim to secure LLMs by addressing vulnerabilities in how they interpret and execute input prompts. Basic approaches include instruction design, where prompts are engineered with explicit safety directives (e.g., “ignore harmful instructions”). However, heavily tuned models may still prioritize conflicting later instructions, and overly restrictive prompts risk stifling legitimate creativity (li-etal-2024-evaluating-instruction; weng2023attack). Similar to pre-processing data augmentation, some techniques attempt to clean the prompt just before the LLM sees it. This can involve paraphrasing prompts or retokenizing inputs to disrupt adversarial patterns, offering limited protection against sophisticated attacks (weng2023attack).
More advanced methods include self-denoising, where LLMs iteratively revise corrupted instructions using their own language understanding to outperform traditional fine-tuning in resisting perturbations (agrawal2025enhancingllmrobustnessperturbed). Further, hu2024large propose RobustGER, which teaches LLMs to perform language-space denoising by incorporating an embedding derived from ASR (Automatic Speech Recognition) N-best lists, refined through audio distillation. Another novel approach, referenced instruction tracking, requires LLMS to explicitly cite which prompt instruction they followed during generation, and so responses referencing malicious injections are automatically filtered, which effectively repurposes the model’s instruction sensitivity as a defence (chen2025robustnessreferencingdefendingprompt). While these intra-processing strategies vary in complexity, they highlight the trade-offs between usability, computational cost, and robustness in real-time LLM safety.
@@
Weighted redistribution refers to adjusting a trained model’s attention weights (without additional training) to reduce bias in its outputs. Since attention weights can reflect learned biases in the data, redistributing them may help the model focus less on biased words or phrases during predictions. pmlr-v162-chai22a developed an adaptive reweighting method that assigns sample-specific weights to prioritize error-prone instances and ensure balanced minority group representation for group-level fairness and exhibits robustness to label noise on various benchmark datasets. Further, considering the fact that AI models focus on specific words (attention weights) may reinforce biases. zhong2025optimizingrlhftraininglarge propose a framework that employs two complementary fusion strategies. First, a data-aware inter-stage fusion overlaps generation and inference by migrating long-tailed samples. Second, a model-aware intra-stage fusion uses a fused pipeline schedule to mitigate pipeline bubbles during training, thereby improving preparation for inference tasks. Here, a pipeline bubble refers to the idle time or inefficiency that occurs in pipeline-parallel training when some stages of the pipeline are waiting for others to complete their work, rather than performing computations. It doesn’t make a model less robust, but inefficient training caused by bubbles can indirectly reduce robustness.
@@
The methods that guide how LLMs select tokens to generate text also play a subtle but important role in balancing output quality and robustness. While standard approaches like greedy decoding (choosing the most probable token) or beam search are widely used, they often struggle under adversarial conditions or when generating diverse, reliable outputs (beyer2025llmsafetyevaluationslackrobustness). Despite their limitations, research has largely prioritized input- or output-level defences (e.g., prompt engineering, response filtering) over reimagining decoding itself as a primary robustness mechanism. This gap persists even as adversarial attacks exploit decoding weaknesses, such as using greedy search or Greedy Coordinate Gradient (GCG) (kumar2025certifyingllmsafetyadversarial) to craft optimal input perturbations. The relative neglect of defensive decoding strategies may stem from practical hurdles, such as the computational complexity of altering core generation processes compared to simpler input/output interventions.
Emerging approaches aim to address these gaps by refining how LLMs navigate token selection. Uncertainty-aware decoding, for instance, could suppress generations when the model’s confidence is low, reducing errors in high-stakes scenarios. Ensemble decoding combines outputs from varied model states or prompts to dilute adversarial influences, while constrained decoding enforces safety rules (e.g., blocking harmful phrases) during token selection (shi2024decodingtime; gu2024charedcharacterwiseensembledecoding; daheim2025uncertaintyaware). son2025robustmultiobjectivecontrolleddecoding formalizes this challenge as a maximin game, introducing RMOD (Robust Multi-Objective Decoding), a method aimed at maximizing worst-case rewards through Nash equilibrium optimization. RMOD achieves this by balancing competing objectives, specifically, identifying a Nash equilibrium between reward weight allocations and the sampling policy during text generation.
@@
Beyond prompt manipulation and decoding, other intra-processing techniques involve adapting the model or its execution environment at inference time. While efficient inference infrastructure does not directly enhance robustness, it enables the deployment of adaptive systems that can improve real-world reliability (manvi2024adaptiveinferencetimecomputellms). For instance, levine2024baselineanalysisrewardmodels evaluates the performance and calibration of reward models under distribution shift and finds that accuracy degrades significantly, particularly for OOD responses. Some of key strategies include test-time adaptation (TTA), which makes minor, instance-specific adjustments to improve handling of rare or unfamiliar data, such as medical cases outside training distributions (snell2024scalingllmtesttimecompute). Architectural interventions explore structural modifications during inference, like removing or swapping transformer layers, revealing that models retain most functionality despite such changes, which is a sign of inherent redundancy and fault tolerance (lad2024remarkablerobustnessllmsstages). Other work examines components like layer normalization, showing how architectural choices impact stability under perturbations (jha2024aerosoftmaxonlyllmsefficient). Finally, efficient inference engines optimize computational bottlenecks (e.g., attention mechanisms) using memory optimization and adaptive computation, enabling complex real-time processes like retrieval-augmented generation while balancing speed and accuracy (ye2025flashinferefficientcustomizableattention). Together, these approaches highlight how real-time adaptability and structural insights can bolster LLM reliability without sacrificing efficiency.